Monochromatic light of frequency 6.0x10^14 Hz is produced by a laser. The power emitted is 2x10^-3 W. The number of photons emitted, on the average, by the source per second is?

Grade:12th pass

2 Answers

Yogesh Saravanan
7 years ago
energy of one photon is = h* freq
power is the energy per unit time 
therfore the energy given per second is 2 * 10^-3
 
number of photons * energy of one photon = total energy 
 
number of photons = (2*10^-3)/(6.636*10^-34 * 6 * 10^14)
 approxi mately = 5 * 10^15 photons . 
Yash Chourasiya
3 years ago
Dear Student

An energy,
E = hv
E= 6.63×10–34 x 6.05×1014
E = 3.98×10–19J
Now, number of photons emitted per second n = p/E
= 2.0 x 10-3/3.98 x 10-19J
= 5 x 1015photons per second.

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya